OpenStudy (anonymous):

Find the sum: \[1+1+\frac{\sin2x}{\sin^2x}+\frac{\sin3x}{\sin^3x}+\cdots+\frac{\sin kx}{\sin^kx}\]

OpenStudy (anonymous):

Hint: DeMoivre's theorem is your friend

ganeshie8 (ganeshie8):

@eliassaab

OpenStudy (anonymous):

*

OpenStudy (xapproachesinfinity):

i wonder what summation this will have

OpenStudy (anonymous):

This series does not seem to be even convergent for every x

OpenStudy (xapproachesinfinity):

what did you do to prove it's divergent?

OpenStudy (xapproachesinfinity):

1+1+sum (k=0 to oo)(sin(nx)/sin^n(x))

OpenStudy (anonymous):

All the terms containing sin are infinite at x =0, or I might be missing something

OpenStudy (xapproachesinfinity):

how is DeMoiver will help

OpenStudy (xapproachesinfinity):

true i appears to me the limit is infinity for those terms?

OpenStudy (xapproachesinfinity):

it*

OpenStudy (anonymous):

The original problem might have been a finite series up to \(k\) terms, I might be remembering it incorrectly.

OpenStudy (anonymous):

Here's one approach. Let \[S_{\text{sine}}=\sum_{n=1}^k\frac{\sin nx}{\sin^nx}=\frac{\sin x}{\sin x}+\frac{\sin2x}{\sin^2x}+\cdots+\frac{\sin kx}{\sin^kx}\] and \[S_{\text{cosine}}=\sum_{n=1}^k \frac{\cos nx}{\sin^nx}=\frac{\cos x}{\sin x}+\frac{\cos2x}{\sin^2x}+\cdots+\frac{\cos kx}{\sin^kx}\] and so \[\begin{align*} S_{\text{cosine}}+iS_{\text{sine}}&=\frac{\cos x+i\sin x}{\sin x}+\frac{\cos2x+i\sin2x}{\sin^2x}+\cdots+\frac{\cos kx+i\sin kx}{\sin^kx}\\ &=\sum_{n=1}^k\frac{e^{inx}}{\sin^nx} \end{align*}\] (Now that I see it worked out again, I believe it was a finite series after all. Edited.) From here it's a matter of isolating the sine series.